OpenStudy (anonymous):

Solve using the quadratic formula x^2 +5×+6=0

OpenStudy (anonymous):

Do you have the formula on you?

OpenStudy (anonymous):

x = [-b +- sqrt(b^2 - 4ac)]/2a

OpenStudy (anonymous):

When a = 1 and b = 5 and c = 6.

OpenStudy (anonymous):

x = [-5 +- sqrt(25 - 24)]/2

OpenStudy (anonymous):

x = [-5 +- sqrt(1)]/2

OpenStudy (anonymous):

@average_chick ?

OpenStudy (anonymous):

x = [-5 - 1]/2 x = [-5 + 1]/2 x = -6/2 x = -4/2